# A Zero-Added-Loss Multiplexing (ZALM) Source Simulation

### v1

@@ Abstract
Zero-Added-Loss Multiplexing (ZALM) offers broadband, per-channel–heralded EPR pairs, with a rich parameter space that allows its performance to be tailored for specific applications. We present a modular ZALM simulator that demonstrates how design choices affect output rate and fidelity. Built in NetSquid with QSI controllers, it exposes 20+ tunable parameters, supports ’IDEAL’ and ’REALISTIC’ modes, and provides reusable components for Spontaneous Parametric Down Conversio (SPDC) sources, interference, Dense Wavelength Division Multiplexing (DWDM) filtering, fiber delay, active polarization gates, detectors, and lossy fiber. Physics-based models capture Hong–Ou–Mandel (HOM) visibility, insertion loss, detector efficiency, gate errors, and attenuation. Using this tool, we map trade-offs among fidelity, link distance, and entangled pairs per use, and show how SPDC bandwidth and DWDM grid spacing steer performance. Using the default configuration settings, average fidelity remains constant at 0.8 but the ebit rate decreases from 0.0175 at the source to 0.0 at 50 km; narrowing the SPDC degeneracy bandwidth increases the ebit rate significantly without affecting fidelity. The simulator enables co-design of source, filtering, and feed-forward settings for specific quantum memories and integrates as a building block for end-to-end quantum-network studies.
@@ I Introduction
Robust entanglement distribution is a foundational requirement for the future quantum internet, as laid out by Kimble [1] and Wehner et al. [2]. While much research has focused on optimising the rate and fidelity of this entanglement distribution, the next step is integrating with quantum memories to form functional quantum memory networks. Adopting the principles of systems engineering, we argue that a holistic, end-to-end analysis of the network architecture is critical. This perspective integrates the performance of quantum entanglement sources, the fidelity degradation across quantum channels, and the efficiencies of quantum memories. To address this, we model a novel quantum source specifically engineered for high-efficiency loading into quantum memories, thereby improving the performance of the entire quantum network. Quantum entanglement sources come in many forms, with different natural wavelengths, emission efficiencies, and bandwidths, as described by Eisaman et al. [3]. Placing these sources into two main categories, we can look at their emission properties as either deterministic, such as the very precise wavelengths emitted by quantum dots, or probabilistic, whereby the output photons will be found within a specific bandwidth centred at a known wavelength. Although deterministic sources simplify network setups, the broadband capabilities of probabilistic sources allow for greater use of the available spectrum and enable the use of common classical network management techniques, such as routing and spectrum allocation, see Bali et al. [4]. To scale quantum networks beyond simple point-to-point links, quantum memories are an essential component. They provide the crucial capability to synchronise operations between distant nodes and to mitigate the inherently probabilistic nature of entanglement generation and swapping protocols. As such, memories form the cornerstone of quantum repeater architectures. The search for a practical physical implementation has led to significant research in solid-state systems. For instance, Weber et al. [5] showed that defects, or vacancies, in diamond are a practical and robust mechanism to provide room-temperature quantum memories. Further work by Bradec et al. [6] highlighted the attributes of the other group IV elemental defects in diamond, with Silicon showing considerable coherence times, 1s, and a more efficient spin-photon interface. However, successfully interfacing with these memories is a significant challenge, with Bersin et al. [7] and Raymer et al. [8] detailing the specific requirements on the incoming photons’ properties, such as bandwidth and wavelength matching. The Zero-Added-Loss Multiplexing (ZALM) protocol introduced by Chen et al. [9] provides for a broadband source with high memory loading efficiency. As there is currently no agreed implementation of the transmitter (Tx) component of this protocol and there are numerous configuration options of each of the transmitter subcomponents, we created a configurable simulation to determine the optimal transmitter architecture for maximising entanglement distribution rates and fidelity.
@@ II The ZALM Source
ZALM, as described by Chen et al. [9] and Shapiro et al. [10], is a broadband, deterministic source-in-the-middle architecture that is optimised for memory loading through the spin-photon interface. ZALM combines a pair of probabilistic SPDC sources into a deterministic quantum entanglement source where the timing and wavelength of a quantum entanglement pair emission is known by the source and communicated to the receiver in advance of transmission. This information is used by the receiving node to efficiently convert the incoming photons for maximum compatibility with the memory loading process, hence there is Zero-Added-Loss at the quantum receiver. Whilst the emission rate of the ZALM source may be lower than other quantum entanglement sources, it has the advantage of only emitting entangled pairs that are capable of being loaded into a memory at the receiver. It classically heralds the receivers to expect a photon within a certain time frame allowing the unheralded (non entangled) photons to be discarded.
The ZALM source consists of (i) two SPDC processes which are triggered simultaneously. These SPDC processes are configured to each produce a two-qubit entangled state. The qubits are physically photons111The terms Photons and Qubits are used interchangeably in the text. When discussing a photon, the authors generally discuss the physical properties associated with the photon and not the quantum state. When qubits are used, the authors focus on the quantum state of the system. that are polarization encoded. The photons produced by each SPDC process are labelled as signal and idler. The signal photons go directly to (ii) a delay line for later processing, whilst the idler photons are passed through (iii) a beam splitter (BS).
The beam splitter, with some probability, will entangle the two idler photons, producing a four-qubit quantum state (which includes the two signal photons in the delay lines). The two idler photons are then emitted from the beam splitter where they pass through (iv) a polarising beam splitter (PBS), which separates them into their respective states. The PBSs are connected to (v) DWDM filters, each of which have as many (vi) detectors attached as channels available. Therefore, each detector has an associated central wavelength for the idler photon with a margin of error < +/- 50% of the DWDM channel width. The coincidence counts at the detectors are then heralded at the (vii) heralding station, which notifies the (viii) post-processor any corrections to be applied to the signal photons.
@@ III The ZALM Source Model
The ZALM source produces entanglement in multiple stages. Initially, entanglement is generated using SPDC processes, where a crystal is placed in a sagnac configuration, see Fig. 2. Two SPDC processes are used in the ZALM source and each creates, with some level of probability, a two-qubit quantum entangled state. These separate quantum states are then entangled, again with some level of probability, with each other to create a four-qubit state. These stages are now described in more detail.
@@ III-A Spontaneous Parametric Down Conversion
There are three types of SPDC process; 0, I, and II. Each SPDC process is pumped by a laser, but the conservation of energy in each process type acts differently, which can affect the wavelength, emission angle, polarisation, and the bandwidth of the emitted photons.
@@ III-B1 Beam Splitter
The beam splitter has 2 input ports and 2 output ports. Each input port is connected to the idler output from an SPDC process, and each output port is connected to the input port of a polarising beam splitter. One of the output ports, the reflected port, adds a 180°phase shift. Photons will both exit at any one port if they are indistinguishable; otherwise there is a 50:50 chance that they will either exit together through just one port or they will exit separate ports. The exit port(s) is a random selection.
As the SPDC process creates an H and V polarised pair of photons, any combination of H and V (HH, HV, VH, VV) can reach the beam splitter resulting in two scenarios when the photons are distinguishable (orthogonal polarisation) and two scenarios whereby they are indistinguishable.
Thus, statistically, in a perfect system, 50% of the time the photons will be indistinguishable and will transmit out a single port and a further 50% of the remaining photons will also transmit out a single port. Therefore, in an ideal system, 75% of the photons will transmit out a single port. Furthermore, the beam splitter may introduce a phase shift, which in the global setting can be ignored, thereby enabling the creation of states and .
@@ III-B2 Hong–Ou–Mandel Visibility
No system is ideal, and a major source of decoherence is introduced by the Hong–Ou–Mandel (HOM) visibility of the photons that interact in the beam splitter, see Hong et al. [11]. In the case of the ZALM source, the HOM visibility is concerned with how closely the two photons overlap spectrally, temporally, and in polarisation, with perfect overlapping in all three degrees of freedom producing completely indistinguishable photons. Anything less than perfect reduces the fidelity of the entanglement to the point that fidelity is zero if the photons are completely distinguishable.
@@ III-B3 Polarising Beam Splitter
If both photons are indistinguishable, they will exit the same port on both the beam splitter and the polarising beam splitter and will therefore create a or Bell state depending on whether they exited the or arm of the beam splitter. Photon Number Resolving detectors will be required to register the double click.
@@
In the final scenario, both photons will exit separate ports on the beam splitter, where one photon will exit the H arm of one of the polarising beam splitters and the second photon the V arm of the other polarising beam splitter. This will produce a Bell state as there is a phase difference between both photons. See Table I for a description of the polarisation encoded Bell states.
@@ III-C DWDM Filtering and Detection
A DWDM filter is connected to each output of the polarising beam splitters. DWDM filters are assumed to have a detector per DWDM channel, regardless of channel bandwidth.
@@ III-C1 DWDM Channels
Four identical DWDM filters are used to determine the approximate wavelength of the idler photons. These DWDM filters will typically use a grid spacing slightly larger than the emitted photon bandwidth and the guardband bandwidth to provide the most accurate wavelength estimate. It is this specific feature that introduces the Zero-Added-Loss characteristics of ZALM.
@@ III-D Post Processing
Based on the Bell state detected, some operations may be applied the signal photons to transform them to the state333This action could be performed at the receiving side.. A fibre delay loop is placed between the signal port on the SPDC process and Pockels cells which are tuned to implement the Pauli gates necessary.
@@ III-D1 Delay Line
The singlet state is the preferred choice for the quantum memory, as it can be loaded with very high (unit) fidelity using the augmented Duan-Kimble protocol described by Raymer et al. [8].
Lastly, Shapiro et al. [10] provide a simple circuit that could be used to convert the polarisation encoded qubits into time-bin encoded qubits for more robust transmission over fibre.
@@ IV Implementation / Results
This ZALM source simulation is implemented in NetSquid, developed by Coopmans et al. [12], and uses the QSI Controller and APIs from Sekavčnik[13]. It can run standalone in NetSquid, or as part of a larger system within QSI.
The ZALM source is modelled on a table-top configuration where each component with conditional output paths is a distinct Node, i.e. the beam splitter is a separate node whereas the DWDM filters and the associated detectors are combined as the detectors have no output paths, rather they only register clicks. Each node has a quantum processor attached to it, which permits quantum operations (gates) to be applied to the qubits. Quantum noise models are applied to these quantum processors, as described in Section IV-A. Furthermore, each node has a set of input and output ports, quantum channels, that are connected together via fibre of configurable length. Lastly, there is a classical information channel between the HeraldingStation and the PostProcessors.
A NodeProtocol runs on each node. The NodeProtocol is specific to the function of the node / component - such as; BeamSplitterProtocol, FilterProtocol, PostProcessorProtocol, etc. - and will execute a QuantumProgram when direct manipulation of qubits is required, i.e. SPDCInitQubits, SPDCProgram, etc. Signals are used to communicate between the NodeProtocols. These Signals act as triggers to other NodeProtocols and are used to pass monitoring / status information.
A FlyingQubit class, inherited from type Qubit, was created (custom_qubits.py) to add additional properties, frequency and polarisation, to the default NetSquid qubit. A density matrix formalism is used to represent the quantum states. All qubits are stored in the quantum memory associated with each node and are popped when an operation needs to be performed, including any gate or depolarising operations, and are then either placed back into memory or transmitted to the next port. The fibres between nodes have a fibre noise model applied.
A utils.py library provides supporting functions to the model, such as calculate_visibility, determineBellState, generate_flex_dwdm_grid, etc., system_setup.py builds and connects the node configuration into a network, and config.py stores the default configuration parameters described in Section IV-A. These parameters can be overridden in the config file or via the QSI API.
@@ IV-A Configuration options
The model can be run in either of two modes; ’IDEAL’ or ’REALISTIC’. The IDEAL mode introduces no losses or noise sources into the model, which is useful for tracking the quantum states through the system, whereas the REALISTIC mode operates with standardised values from real-world components.
| SIM_MODE | ’REALISTIC’ |
| ’IDEAL’ |
The default values for each configuration option in REALISTIC mode are now described.
The SPDC process has 6 configurable parameters which specify the pump laser wavelength (nm), the degeneracy bandwidth of the SPDC processes (nm), the jitter between both processes (ps), the output photon bandwidth (GHz), the probability of a successful SPDC emission, and the signal/idler separation method.
@@
| TEMPORAL_JITTER_STDEV_PS | 20.0 |
@@
| SPDC_MODE | ’DICHROIC’ |
| ’PBS’ |
The Beam Splitter has 2 configurable parameters which specify the minimum HOM visibility to be determined indistinguishable and the insertion loss associated with it.
@@
The Polarising Beam Splitter has 2 configurable parameters which specify the probability of a photon being emitted from the incorrect port and the insertion loss associated with it.
@@
The DWDM Filter has 6 configurable parameters which specify which DWDM bands are enabled, the DWDM grid bandwidth (GHz), the size of any guardbands, the insertion loss associated with it, and the filtering model in use.
@@
| FILTER_MODEL | ’GAUSSIAN’ |
| ’BRICKWALL’ |
The Detectors have 2 configurable parameters, their detection efficiency and whether they are photon number resolving.
@@
| DETECTOR_TYPE | ’STANDARD |
| ’PNR’ |
The QuantumProcessors have 4 configurable parameters, the gate error probability over a single gate, the gate error probability over two gates, the measurement dephasing probability, and the memory depolarization rate.
@@
The fibre INTERNODE_LENGTH (km) is configurable between the heralding station and the measurement station to simulate a quantum channel.
| INTERNODE_LENGTH | 15 |

### v3

@@ Abstract
Zero-Added-Loss Multiplexing (ZALM) offers broadband, per-channel–heralded EPR pairs, with a rich parameter space that allows its performance to be tailored for specific applications. We present a modular ZALM simulator that demonstrates how design choices affect output rate and fidelity. Built in NetSquid with QSI controllers, it exposes 20+ tunable parameters, supports ’IDEAL’ and ’REALISTIC’ modes, and provides reusable components for Spontaneous Parametric Down Conversion (SPDC) sources, interference, Dense Wavelength Division Multiplexing (DWDM) filtering, fiber delay, active polarization gates, detectors, and lossy fiber. Physics-based models capture Hong–Ou–Mandel (HOM) visibility, insertion loss, detector efficiency, gate errors, and attenuation. Using this tool, we map trade-offs among fidelity, link distance, and entangled pairs per use, and show how SPDC bandwidth and DWDM grid spacing steer performance. Using the default configuration settings, average fidelity remains constant at 0.83 but the ebit rate decreases from 0.0175 at the source to 0.0 at 50 km; narrowing the SPDC degeneracy bandwidth increases the ebit rate significantly without affecting fidelity. The simulator enables co-design of source, filtering, and feed-forward settings for specific quantum memories and integrates as a building block for end-to-end quantum-network studies.
@@ I Introduction
Robust entanglement distribution is a foundational requirement for the future quantum internet, as laid out by [kimble2008quantum] and [wehner2018quantum]. While much research has focused on optimising the rate and fidelity of this entanglement distribution, the next step is integrating with quantum memories to form functional quantum memory networks. Adopting the principles of systems engineering, we argue that a holistic, end-to-end analysis of the network architecture is critical. This perspective integrates the performance of quantum entanglement sources, the fidelity degradation across quantum channels, and the efficiencies of quantum memories. To address this, we model a novel quantum source specifically engineered for high-efficiency loading into quantum memories, thereby improving the performance of the entire quantum network.
Quantum entanglement sources come in many forms, with different natural wavelengths, emission efficiencies, and bandwidths, as described by [eisaman2011invited]. Placing these sources into two main categories, we can look at their emission properties as either deterministic, such as the very precise wavelengths emitted by quantum dots, or probabilistic, whereby the specific emission wavelength of a generated photon pair is a stochastic variable distributed across a broad spectral bandwidth centred at a known degeneracy point. Although deterministic sources simplify network setups, the broadband capabilities of probabilistic sources allow for greater use of the available spectrum. This is enabled by adapting common classical network management techniques for routing and spectrum allocation, see [jsac_rsa].
To scale quantum networks beyond simple point-to-point links, quantum memories are an essential component. They provide the crucial capability to synchronise operations between distant nodes and to mitigate the inherently probabilistic nature of entanglement generation and swapping protocols. As such, memories form the cornerstone of quantum repeater architectures. The search for a practical physical implementation has led to significant research in solid-state systems. For instance, [weber2010quantum] showed that defects, or vacancies, in diamond are a practical and robust mechanism to provide quantum memories. Further work by [bradac2019quantum] highlighted the attributes of the other group IV elemental defects in diamond, with Silicon showing considerable coherence times, > 1s, and a more efficient spin-photon interface. However, successfully interfacing with these memories is a significant challenge, with [bersin2024telecom] and [PhysRevApplied.22.044013] detailing the specific requirements on the incoming photons’ properties, such as bandwidth and wavelength matching.
The Zero-Added-Loss Multiplexing (ZALM) protocol introduced by [Chen_ZALM] provides for a broadband source with high memory loading efficiency. As there is currently no agreed implementation of the transmitter (Tx) component of this protocol and there are numerous configuration options of each of the transmitter subcomponents, we created a configurable simulation to determine the optimal transmitter architecture for maximising entanglement distribution rates and fidelity.
@@ II The ZALM Source
ZALM, as described by [Chen_ZALM] and [shapiro2024entanglement], is a broadband, deterministic source-in-the-middle architecture that is optimised for memory loading through the spin-photon interface. ZALM combines a pair of probabilistic SPDC sources into a quasi-deterministic quantum entanglement source where an entangled pair is emitted randomly but its timing and emission wavelength is known by the source and communicated to the receiver in advance of transmission. This information is used by the receiving node to efficiently convert the incoming photons for maximum compatibility with the memory loading process, hence there is ‘zero added loss’ at the quantum receiver.
Whilst the emission rate of the ZALM source may be lower than other quantum entanglement sources, it has the advantage of only emitting entangled pairs that are capable of being loaded into a memory at the receiver. It classically heralds the receivers to expect a photon within a certain time frame allowing the unheralded (non entangled) photons to be discarded.
The ZALM source consists of (i) two SPDC processes which are triggered simultaneously. These SPDC processes are configured to each produce a two-qubit entangled state (ebit). The qubits are physically photons111The terms ‘photons’ and ‘qubits’ are used interchangeably in the text. When discussing a photon, the authors generally discuss the physical properties associated with the photon and not the quantum state. When qubits are used, the authors focus on the quantum state of the system. that are polarisation encoded. The photons produced by each SPDC process are labelled as signal and idler. The signal photons go directly to (ii) a delay line for later processing, whilst the idler photons are passed through (iii) a beam splitter (BS).
The beam splitter, with some probability, will entangle the two idler photons, producing a four-qubit quantum state (which includes the two signal photons in the delay lines). The two idler photons are then emitted from the beam splitter where they pass through (iv) a polarising beam splitter (PBS), which separates them into their respective states. The PBSs are connected to (v) DWDM filters, each of which have as many (vi) detectors attached as channels available. Therefore, each detector has an associated central wavelength for the idler photon with a margin of error within of the DWDM channel width. The coincidence counts at the detectors are then heralded at the (vii) heralding station, which notifies the (viii) post-processor any corrections to be applied to the signal photons.
@@ III The ZALM Source Model
The ZALM source produces entanglement in multiple stages. Initially, entanglement is generated using SPDC processes, where a crystal is placed in a Sagnac configuration, see Fig. 2. Two SPDC processes are used in the ZALM source and each creates, with some probability, a two-qubit quantum entangled state. These separate quantum states are then entangled, again with some probability, with each other to create a four-qubit state. These stages are now described in more detail.
@@ III-A Spontaneous Parametric Down Conversion
There are three types of SPDC processes: 0, I, and II. Each SPDC process is pumped by a laser, but the conservation of energy in each type acts differently, which affects the wavelength, emission angle, polarisation, and the bandwidth of the emitted photons.
@@ III-B1 Beam Splitter
The beam splitter has 2 input ports and 2 output ports. Each input port is connected to the idler output from an SPDC process, and each output port is connected to the input port of a polarising beam splitter. One of the output ports, the reflected port, adds a 180° phase shift. Photons will both exit at any one port if they are indistinguishable; otherwise there is a 50:50 chance that they will exit together through a single port or separate ports. The exit port(s) are selected randomly.
As the SPDC process creates an H and V polarised pair of photons, any combination of H and V (HH, HV, VH, VV) can reach the beam splitter. This results in two scenarios when the photons are distinguishable (orthogonal polarisation) and two scenarios whereby they are indistinguishable.
Thus, statistically, in an ideal system, 50% of the time the photons will be indistinguishable and will exit a single port and a further 50% of the remaining photons will also exit a single port. Therefore, in an ideal system, 75% of the photons will exit a single port. Furthermore, the beam splitter may introduce a phase shift, which in the global setting can be ignored, thereby enabling the creation of states and .
@@ III-B2 Hong–Ou–Mandel Visibility
No system is ideal, and a major source of decoherence is introduced by the Hong–Ou–Mandel (HOM) visibility of the photons that interact in the beam splitter, see [HOMvisibility]. In the case of the ZALM source, the HOM visibility is concerned with how closely the two photons overlap spectrally, temporally, and in polarisation, with perfect overlapping in all three degrees of freedom producing completely indistinguishable photons. Anything less than perfect reduces the fidelity of the entanglement to the point that fidelity is zero if the photons are completely distinguishable.
@@ III-B3 Polarising Beam Splitter
If both photons are indistinguishable, they will exit the same port on both the beam splitter and the polarising beam splitter and will therefore create a or Bell state depending on whether they exited the or arm of the beam splitter. Photon-number-resolving detectors will be required to register the double click.
@@
In the final scenario, both photons will exit separate ports on the beam splitter, where one photon will exit the H arm of one of the polarising beam splitters and the second photon the V arm of the other polarising beam splitter. This will produce a Bell state as there is a phase difference between both photons. See Table I for a description of the polarisation-encoded Bell states.
@@ III-C DWDM Filtering and Detection
A DWDM filter is connected to each output of the polarising beam splitters. DWDM filters are assumed to have one detector per DWDM channel, regardless of channel bandwidth.
@@ III-C1 DWDM Channels
Four identical DWDM filters are used to determine the approximate centre wavelength of the idler photons. These DWDM filters will typically use a grid spacing slightly larger than the emitted photon bandwidth and a guard-band to estimate the centre wavelength accurately. It is this specific feature that introduces the zero added loss characteristics of ZALM.
@@ III-D Post Processing
Based on the Bell state detected, some operations may be applied to the signal photons to transform them to the state333This action could be performed at the receiving side.. A fibre delay loop is placed between the signal port on the SPDC process and Pockels cells which are tuned to implement the Pauli gates necessary.
@@ III-D1 Delay Line
The singlet state is the preferred choice for the quantum memory, as it can be loaded with very high (unit) fidelity using the augmented Duan-Kimble protocol described by [PhysRevApplied.22.044013].
Lastly, [shapiro2024entanglement] provide a simple circuit that could be used to convert the polarisation encoded qubits into time-bin encoded qubits for more robust transmission over fibre.
@@ IV Implementation / Results
This ZALM source simulation is implemented in NetSquid, developed by [coopmans2021netsquid], and uses the QSI Controller and APIs from [QSI_Library_2022]. It can run standalone in NetSquid, or as part of a larger system within QSI.
The ZALM source is modelled on a table-top configuration where each component with conditional output paths is a distinct Node, i.e. the beam splitter is a separate node whereas the DWDM filters and the associated detectors are combined as the detectors have no output paths, rather they only register clicks. Each node has a quantum processor attached to it, which permits quantum operations (gates) to be applied to the qubits, Fig. 5 details the quantum operations (gates/measurements) in a QisKit-style quantum circuit. Quantum noise models are applied to these quantum processors, as described in Section IV-A. Furthermore, each node has a set of input and output ports, quantum channels, that are connected together via fibre of configurable length. Lastly, there is a classical information channel between the HeraldingStation and the PostProcessors.
A NodeProtocol runs on each node. The NodeProtocol is specific to the function of the node / component – such as; BeamSplitterProtocol, FilterProtocol, PostProcessorProtocol, etc. – and will execute a QuantumProgram when direct manipulation of qubits is required, i.e. SPDCInitQubits, SPDCProgram, etc. Signals are used to communicate between the NodeProtocols. These Signals act as triggers to other NodeProtocols and are used to pass monitoring / status information. The schematic in Fig. 4 shows the ZALMProtocol, its sub-protocols, which can be mapped directly to the physical components in Fig. 1, and the signalling between them.
A FlyingQubit class, inherited from type Qubit, was created (custom_qubits.py) to add additional properties, frequency and polarisation, to the default NetSquid qubit. A density matrix formalism is used to represent the quantum states. All qubits are stored in the quantum memory associated with each node. They are popped when an operation needs to be performed, including any gate or depolarising operations, and are then either placed back into memory or transmitted to the next port. The fibres between nodes have a fibre noise model applied.
A utils.py library provides supporting functions to the model, such as calculate_visibility, determineBellState, generate_flex_dwdm_grid, etc., system_setup.py builds and connects the node configuration into a network, and config.py stores the default configuration parameters described in Section IV-A. These parameters can be overridden in the config file. Additional implementation details are provided in Appendices B and C.
@@ IV-A Configuration options
The model can be run in either of two modes by setting ‘SIM_MODE’ to either ‘IDEAL’ or ‘REALISTIC’. The IDEAL mode introduces no losses or noise sources into the model, which is useful for tracking the quantum states through the system, whereas the REALISTIC mode operates with standardised values from real-world components. We now describe the default values for each configuration option in REALISTIC mode.
The SPDC process has 6 configurable parameters which specify the pump laser wavelength (nm), the degeneracy bandwidth of the SPDC processes (nm), the jitter between both processes (ps), the output photon bandwidth (GHz), the probability of a successful SPDC emission, and the signal/idler separation method:
@@
| TEMPORAL_JITTER_STDEV_PS | 0.0 |
@@
| SPDC_MODE | ‘DICHROIC’ |
| ‘PBS’ |
The BeamSplitter has 2 configurable parameters which specify the minimum HOM visibility to be determined indistinguishable and the insertion loss associated with it:
@@
The Polarising Beam Splitter has 2 configurable parameters which specify the probability of a photon being emitted from the incorrect port and the insertion loss associated with it:
@@
The DWDM Filter has 6 configurable parameters which specify which DWDM bands are enabled, the DWDM grid bandwidth (GHz), the size of any guard-bands, the insertion loss associated with it, and the filtering model in use:
@@
| FILTER_MODEL | ‘GAUSSIAN’ |
| ‘BRICKWALL’ |
The Detectors have 2 configurable parameters, their detection efficiency and whether they are photon number resolving:
@@
| DETECTOR_TYPE | ‘STANDARD’ |
| ‘PNR’ |
The QuantumProcessors have 4 configurable parameters, the gate error probability over a single gate, the gate error probability over two gates, the measurement dephasing probability, and the memory depolarization rate:
@@
The fibre INTERNODE_LENGTH (km) is configurable between the heralding station and the measurement station to simulate a quantum channel:
| INTERNODE_LENGTH | 1 |